# Quantifying the Balinski-Young Theorem: Structure and Probability of Quota Violations in Divisor Methods for Three States

### v2

@@ 5.3.1. Motivating Example
5.3.2. Calculation of the Feasible Region
@@
  1. (1)
    (Subset of the quota feasible region) For Dean’s, Adam’s, or Huntington-Hill’s method, and for all pairs ($\lfloor q_{1}\rfloor$$\lfloor q_{2}\rfloor$), $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)\subseteq\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$. Indeed if a point in $\bigtriangleup_{l}$ lay outside this set, a lower quota violation would occur for a state that does not have the largest population, contradicting Lemma 5.3.
  2. (2)
@@
    where then $M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor=\lfloor q_{3}\rfloor+1$. By symmetry, the same holds for the $x$-intercept. In particular, this holds in Adam’s method, Dean’s method, and the Huntington-Hill Method.
@@
5.3.3. Probability Formula for Lower Quota Violations
@@
Let $(q_{1},q_{2})$ be chosen uniformly from $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$. Then, for Dean’s, Adams’s, and Huntington-Hill’s method with $M>3$, the probability of a quota violation is:
@@
5.3.4. Empirical Results
Given an apportionment method $A$, $M$ seats, and $n=3$ states, the probability of a lower quota violation under a uniform distribution is computed using Corollary 5.9.1. The resulting probabilities are shown in the tables below and compared with sample probabilities from 100,000 random samples of $(q_{1},q_{2})$ from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
@@ 5.3.4. Empirical Results
Quota Violations in Adams’s Method with Uniform Quotas
@@
Quota Violations in Dean’s Method with Uniform Quotas
@@
6. Structure and Probability of Upper Quota Violations in Jefferson’s Method for Three States
6.1. Analogous Theoretical Results
We now provide a similar study of upper quota violations in Jefferson’s method when apportioning three states. Webster’s, Adams’s, Dean’s, and the Huntington-Hill Method have no upper quota violations by Lemma 5.1 so only Jefferson’s method needs to be considered.
@@ 6.1. Analogous Theoretical Results
In Jefferson’s method and $n=3$ states, upper quota violations can only occur on the state with the largest population.
@@
(Classification of Upper Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is Jefferson’s method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be an upper quota violation and the upper quota violation must occur on the third state.
6.1.1. Upper Quota Violation Criteria Test for 3 States
@@
(Upper Quota Violation Criteria Test) Let $A$ be Jefferson’s method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has an upper quota violation if and only if all three of the following statements are true:
  1. (1)
@@
  3. (3)
    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{1}\in\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
@@
From there, in analogy with Theorem 5.7, derive the following probability result:
@@
The proof is omitted as it is similar to the proof of Theorem 5.7. Similar properties of Remark 5.8 apply to this region.
6.1.2. Probability Formula for an Upper Quota Violation
@@
(Probability of Upper Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$. For $q_{1},q_{2}<q_{3}$ and Jefferson’s method, the probability of a upper quota violation is given by:
@@
If $(q_{1},q_{2})$ are chosen uniformly on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$, then for Jefferson’s method the probability of an quota violation for any $M$ is given by:
@@
6.1.3. Empirical Results
Given Jefferson’s method, $M$ seats, and $n=3$ states, the probability of an upper quota violation under a uniform distribution is computed using Corollary 6.7.1. The resulting probabilities are shown in the tables below and compared with sample probabilities from 100,000 random samples of $(q_{1},q_{2})$ from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
Quota Violations in Jefferson’s Method with Uniform Quotas
@@ 6.1.3. Empirical Results
7. Conclusion and Next Steps
This work provides a comprehensive probabilistic analysis of quota violations in divisor methods for three states, advancing the quantitative understanding of the Balinski–Young Theorem. By classifying the structure of both lower and upper quota violations, deriving analytic criteria tests, and developing exact probability formulas, we obtain an understanding how often these apportionment violations occur and under what conditions. The combination of theoretical results with empirical validation provides a foundation for understanding the nature of violations where both the number of states and seats are increased.
The results are significant because they illustrate how probabilistic and geometric techniques are combined to analyze apportionment. Specifically, these methods both quantify the frequency and illuminate the structure of quota violations in divisor methods that are foundational in political science and economics. The ability to move beyond impossibility theorems toward concrete probability calculations represents a step forward in the applied probability analysis of voting and apportionment rules.
Several natural directions for future research emerge from this study. First, the methods developed here invite generalization to larger numbers of states and all divisor methods, where the geometry of quota violations becomes richer and potentially more intricate. Second, extending the framework beyond divisor methods to rank-index methods or even to the entire class of apportionment methods covered by the Balinski–Young Theorem would provide a unified probabilistic perspective on fairness violations in apportionment. Third, exploring alternative population distributions, including those motivated by empirical census data, would connect the theoretical findings more closely to the applied settings. Finally, a computational analysis where the number of seats or states grows large could offer insights into the limiting behavior of quota violation probabilities.
@@ 7. Conclusion and Next Steps
8. Declarations
8.1. Ethical Approval and Consent to Participate
@@ 8.2. Competing Interests
8.3. Data Availability Statement

### v3

@@ 5.4.1 Motivating Example
Figure 3: Example Inequality Region without Overlap
5.4.2 Calculation of the Feasible Region
@@
  1. 1.
    (Subset of the quota feasible region) For the Dean, Adam, or Huntington-Hill methods, and for all pairs ($\lfloor q_{1}\rfloor$$\lfloor q_{2}\rfloor$), $\bigtriangleup_{l}(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor)\subseteq\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$. Indeed if a point in $\bigtriangleup_{l}$ lay outside this set, a lower quota violation would occur for a state that does not have the largest population, contradicting Lemma 5.3.
  2. 2.
@@
    where then $M-\lfloor q_{1}\rfloor-\lfloor q_{2}\rfloor=\lfloor q_{3}\rfloor+1$. By symmetry, the same holds for the $x$-intercept. In particular, this holds in Adam’s method, Dean method, and the Huntington-Hill Method.
@@
5.4.3 Probability Formula for Lower Quota Violations
@@
Let $(q_{1},q_{2})$ be chosen uniformly from $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$. Then, for the Dean, Adams, and Huntington-Hill method with $M>3$, the probability of a quota violation is:
@@
5.4.4 Empirical Results
Given an apportionment method $A$, $M$ seats, and $n=3$ states, the probability of a lower quota violation under a uniform distribution is computed using Corollary 5.10.1. The resulting probabilities are shown in the tables below and compared with a simulation of 100,000 random $(q_{1},q_{2})$ samples from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
@@ 5.4.4 Empirical Results
Quota Violations in the Adams Method with Uniform Quotas
@@
Quota Violations in the Dean Method with Uniform Quotas
@@
Comparisons to simulations under different distributions may be found in Appendix A.
6 Structure and Probability of Upper Quota Violations in the Jefferson Method for Three States
6.1 Analogous Theoretical Results
We now provide a similar study of upper quota violations in the Jefferson method when apportioning three states. The Webster, Adams, Dean, and the Huntington-Hill methods have no upper quota violations by Lemma 5.1 so only the Jefferson method needs to be considered.
@@ 6.1 Analogous Theoretical Results
In the Jefferson method and $n=3$ states, upper quota violations can only occur on the state with the largest population.
@@
(Classification of Upper Quota Violations in 3 States) Order the populations $p_{1}<p_{2}<p_{3}$. If $A$ is the Jefferson method and $A(p_{1},p_{2},p_{3})$ has a quota violation, the final apportionment must be $(a_{1},a_{2},a_{3})=(\lfloor q_{1}\rfloor,\lfloor q_{2}\rfloor,\lceil q_{3}\rceil+1)$, where $q_{i}$ is the standard quota of state $i$. That is, the quota violation must be an upper quota violation and the upper quota violation must occur on the third state.
6.1.1 Upper Quota Violation Criteria Test for 3 States
@@
(Upper Quota Violation Criteria Test) Let $A$ be the Jefferson method, with divisor function $d(s)$. Order the populations such that $p_{1}<p_{2}<p_{3}$ and calculate the standard quotas $q_{i}$. Then the apportionment $A(p_{1},p_{2},p_{3})$ on $M$ seats has an upper quota violation if and only if all three of the following statements are true:
  1. 1.
@@
  3. 3.
    $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lceil q_{3}\rceil=M-1$ (equivalently for $q_{3}\notin\mathbb{Z}$, $\lfloor q_{1}\rfloor+\lfloor q_{2}\rfloor+\lfloor q_{3}\rfloor=M-2$.)
@@
From there, in analogy with Theorem 5.8, derive the following result:
@@
The proof is omitted as it is similar to the proof of Theorem 5.8. Similar properties to Remark 5.9 apply to this region.
6.1.2 Geometric Visualization of Upper Quota Violations
Similarly to Section 5.3, one may use Theorem 6.5 and specifically Theorem 6.6 to display all violatory $(q_{1},q_{2},q_{3})$ on the simplex $\{(q_{1},q_{2},q_{3})\mid q_{1}+q_{2}+q_{3}=M\}$. This allows for a visualization of the geometry and structure of upper quota violations. For instance, below is a simplex plot indicating all violatory $(q_{1},q_{2},q_{3})$ in the Jefferson method with $M=13$:
Figure 4: Quota Violations in the Jefferson Method on $\{(q_{1},q_{2},q_{3})\mid q_{1}+q_{2}+q_{3}=M\}$
6.1.3 Probability Formula for Upper Quota Violations
@@
(Probability of Upper Quota Violation) For $M$ seats and $n=3$ states, let $q_{1}$ and $q_{2}$ be the standard quotas of states 1 and 2 with joint pdf $f$, and define $q_{3}=M-q_{1}-q_{2}$. For $q_{1},q_{2}<q_{3}$ and the Jefferson method, the probability of a upper quota violation is given by:
@@
If $(q_{1},q_{2})$ are chosen uniformly on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$, then for the Jefferson method the probability of an quota violation for any $M$ is given by:
@@
6.1.4 Empirical Results
Given the Jefferson method, $M$ seats, and $n=3$ states, the probability of an upper quota violation under a uniform distribution is computed using Corollary 6.7.1. The resulting probabilities are shown in the tables below and compared with sample probabilities from 100,000 random samples of $(q_{1},q_{2})$ from a uniform distribution on $\{(q_{1},q_{2})\mid 0<q_{1}<M-q_{1}-q_{2},0<q_{2}<M-q_{1}-q_{2}\}$:
Quota Violations in the Jefferson Method with Uniform Quotas
@@ 6.1.4 Empirical Results
Comparisons to simulations under different distributions may be found in Appendix A.
7 Conclusion and Next Steps
This work provides a comprehensive analysis of quota violations in divisor methods for three states, advancing the quantitative and qualitative understanding of the Balinski–Young Theorem. The main contribution of the paper lies in making the consequences of the Balinski–Young theorem explicit and quantitative in this restricted setting by providing a complete classification of possible quota violations and analyzing the conditions and frequencies under which they occur. The work thus refines and illustrates a classical impossibility result rather than challenging it or proposing an alternative apportionment method
The results are significant because they illustrate how probabilistic and geometric techniques are combined to analyze apportionment. Specifically, these methods both quantify the frequency and illuminate the structure of quota violations in divisor methods. The ability to move beyond impossibility theorems toward concrete probability calculations represents a step forward in the applied probability analysis of voting and apportionment rules.
Several natural directions for future research emerge from this study. First and foremost, a strong and in progress next step of this work is the generalization to larger numbers of states and all divisor methods, where the geometry of quota violations becomes richer and more intricate. Second, extending the framework beyond divisor methods to rank-index methods or even to the entire class of apportionment methods covered by the Balinski–Young Theorem would provide a unified probabilistic perspective on fairness violations in apportionment. Finally, a computational analysis where the number of seats, states or ratios between populations grows large could offer insights into the limiting behavior of quota violation probabilities.
@@ 7 Conclusion and Next Steps
8 Declarations
8.1 Ethical Approval and Consent to Participate
@@ 8.2 Competing Interests
8.3 Data Availability Statement
@@ 8.5 Funding
Appendix A Comparison of Theoretical Results to Simulations Under Various Distributions
The following are the results of additional simulations and comparisons to theoretical values calculated using Theorems 5.10 and 6.7.
Probability of Quota Violations with $(\frac{q_{1}}{M},\frac{q_{2}}{M},\frac{q_{3}}{M})\sim\mathrm{Dir}(\frac{1}{2},\frac{1}{2},\frac{1}{2})$
Method $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
Huntington-Hill 33 $0.55051$ $0.5527$ $(0.54962,0.55578)$
Adams 33 $0.55051$ $0.55421$ $(0.55113,0.55729)$
Dean 33 $0.55051$ $0.54921$ $(0.54613,0.55229)$
Jefferson 33 $0.01252$ $0.01283$ $(0.01213,0.01353)$
Huntington-Hill 55 $0.33758$ $0.33468$ $(0.33176,0.33760)$
Adams 55 $0.37385$ $0.37233$ $(0.36933,0.37533)$
Dean 55 $0.34624$ $0.34727$ $(0.34432,0.35022)$
Jefferson 55 $0.03003$ $0.02934$ $(0.02829,0.03039)$
Huntington-Hill $10$ $0.20313$ $0.20478$ $(0.20228,0.20728)$
Adams $10$ $0.29367$ $0.29473$ $(0.29190,0.29756)$
Dean $10$ $0.21604$ $0.21577$ $(0.21322,0.21831)$
Jefferson $10$ $0.05637$ $0.05682$ $(0.05539,0.05825)$
Huntington-Hill $15$ $0.15545$ $0.15436$ $(0.15212,0.15660)$
Adams $15$ $0.26701$ $0.26744$ $(0.26470,0.27018)$
Dean $15$ $0.16703$ $0.1651$ $(0.16280,0.16740)$
Jefferson $15$ $0.07108$ $0.07055$ $(0.06896,0.07214)$
Probability of Quota Violations with $(\frac{q_{1}}{M},\frac{q_{2}}{M},\frac{q_{3}}{M})\sim\mathrm{Dir}(2,2,2)$
Method $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
Huntington-Hill 33 $0.13580$ $0.13519$ $(0.13307,0.13731)$
Adams 33 $0.13580$ $0.13486$ $(0.13274,0.13698)$
Dean 33 $0.13580$ $0.13587$ $(0.13375,0.13799)$
Jefferson 33 $0.02670$ $0.02627$ $(0.02528,0.02726)$
Huntington-Hill 55 $0.02276$ $0.2327$ $(0.02234,0.02420)$
Adams 55 $0.03660$ $0.03729$ $(0.03612,0.03846)$
Dean 55 $0.02513$ $0.02497$ $(0.02400,0.02594)$
Jefferson 55 $0.02889$ $0.02862$ $(0.02759,0.02965)$
Huntington-Hill $10$ $0.00386$ $0.00391$ $(0.00352,0.00430)$
Adams $10$ $0.03091$ $0.0304$ $(0.02934,0.03146)$
Dean $10$ $0.00537$ $0.0052$ $(0.00475,0.00565)$
Jefferson $10$ $0.03045$ $0.03104$ $(0.02997,0.03211)$
Huntington-Hill $15$ $0.00150$ $0.00154$ $(0.00130,0.00178)$
Adams $15$ $0.03107$ $0.03195$ $(0.03086,0.03304)$
Dean $15$ $0.00227$ $0.00242$ $(0.00212,0.00272)$
Jefferson $15$ $0.03094$ $0.03015$ $(0.02909,0.03121)$
Remark A.1.
As a Dirichlet $(1,1,1)$ distributions is a uniform distribution on the simplex $\{(x_{1},x_{2},x_{3})\mid x_{1}+x_{2}+x_{3}=1\}$, the case $(\frac{q_{1}}{M},\frac{q_{2}}{M},\frac{q_{3}}{M})\sim\mathrm{Dir}(1,1,1)$ is equivalent to the case presented in Sections 5.4.4 and 6.1.4.
Probability of Quota Violations with $p_{i}$ IID and $p_{i}\sim\exp(1)$
Method $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
Huntington-Hill 33 $0.33333$ $0.33413$ $(0.33121,0.33705)$
Adams 33 $0.33333$ $0.33298$ $(0.33006,0.33590)$
Dean 33 $0.33333$ $0.33379$ $(0.33087,0.33671)$
Jefferson 33 $0.02222$ $0.02199$ $(0.02108,0.02290)$
Huntington-Hill 55 $0.13092$ $0.13225$ $(0.13015,0.13434)$
Adams 55 $0.16000$ $0.15939$ $(0.15712,0.16166)$
Dean 55 $0.13714$ $0.13545$ $(0.13333,0.13757)$
Jefferson 55 $0.03943$ $0.03942$ $(0.03821,0.04063)$
Huntington-Hill $10$ $0.04904$ $0.04949$ $(0.04815,0.05083)$
Adams $10$ $0.11057$ $0.11116$ $(0.10921,0.11311)$
Dean $10$ $0.05581$ $0.05663$ $(0.05520,0.05806)$
Jefferson $10$ $0.05671$ $0.05642$ $(0.05499,0.05785)$
Remark A.2.
The probabilities in the case $p_{i}$ are independent and identically distributed random variables with $p_{i}\sim\exp(\lambda)$ are the same as the probabilities presented in sections 5.4.4 and 6.1.4. This may be shown by computing the associated probability density function for $(q_{1},q_{2})$.
Probability of Quota Violations with $p_{i}$ IID and $p_{i}\sim U(0,1000)$
Method $M$ Theoretical Probability Sample Probability $95\%$ Confidence Interval
Huntington-Hill 33 $0.125$ $0.12422$ $(0.122218,0.12626)$
Adams 33 $0.125$ $0.12399$ $(0.12195,0.12603)$
Dean 33 $0.125$ $0.1251$ $(0.12305,0.12715)$
Jefferson 33 $0.01389$ $0.01341$ $(0.01270,0.1412)$
Huntington-Hill 55 $0.03923$ $0.03966$ $(0.03845,0.04087)$
Adams 55 $0.05903$ $0.05944$ $(0.05797,0.06091)$
Dean 55 $0.04360$ $0.04444$ $(0.04316,0.04572)$
Jefferson 55 $0.01819$ $0.01802$ $(0.01720,0.01884)$
Huntington-Hill $10$ $0.01363$ $0.01357$ $(0.01285,0.01429)$
Adams $10$ $0.04015$ $0.03966$ $(0.03845,0.04087)$
Dean $10$ $0.01612$ $0.01571$ $(0.01494,0.01648)$
Jefferson $10$ $0.02308$ $0.02278$ $(0.02186,0.02370)$
Huntington-Hill $15$ $0.00802$ $0.00807$ $(0.00752,0.00862)$
Adams $15$ $0.03664$ $0.03724$ $(0.03607,0.03841)$
Dean $15$ $0.00953$ $0.01006$ $(0.00944,0.01068)$
Jefferson $15$ $0.02536$ $0.02561$ $(0.02463,0.02659)$